Tech Staff Meeting Notes August 24, 2009

The top priority is always repair of down stations.
NM problem stations
ET problem stations

The task list updated February 6, 2009 needs to be updated again. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


Priorities

Repair down stations
Weigh Station/Mud Island/Boswell project
AG Network design
CVVA trigger upgrades
SM sample rate fix/yr 3 hardware upgrade


News & Highlights

Jim is back but will need to excersize his fingers more than his back for a while.

The CERI welcome back lunch will be at noon on thursday August 27. Staff are encouraged to participate and will be asked to introduce themselves.

Chris is arranging annual safety training for himself, David, and Jim. All others are welcome to attend.

We are approved for the following in approximately the same order as we hope to complete them:

  1. Upgrade 11 NM broadband stations with new 6-c DAS, new sensor, and accelerometer.
  2. Upgrade 2 ET broadband stations with new 6-c DAS, new sensor, and accelerometer.
  3. Upgrade 3 ET shortperiod analog stations with new 6-c das, digital telemetry, and accelerometer.

NMSZ

Short Period & Broadbands

Its likely that some of these station problems are at the nodes and can be fixed by checking the discriminators in detail.

Holly's NM list:
Down
PPLM

Isis Oxidation problem
SJBM.EHN
JHAR.EHE
WYBT.EHZ - and drifting
MARM.EHN

Gain Channel problem
MORT.EHG
CHRM.EHG

Muted Amplitude
GOBM.EHN
MCAM.EHN
LRAR.EHE

Unusable
CHRM

With phone support from Jim, David took care of most of the problems. The canister from PPLM is repaired and ready to reinstall.

Several of the stations can get improved telemetry by transmitting directly to the node. The original configuration changed enough so that there is room in the band. Jim will investigate.

NODE Tasks

MKTA:

LNXT: Still need to tie generator status alarms into lnxtworm

NMAD:

AZTN: Upgrade PC and replace XLAK clock

CERI:

DSL is installed at lnxt and bosewell. Steve needs to complete the config at lnxt. Boswell needs the hardware installed. Telecom said we're too far from the switch to get it at mkta and nmad--we had ISDN at those sites so Steve is checking directly with local provider and, failing that, will investigate other options (e.g. satellite).

Need to gather up a contact list (e.g. everyone's home phone, power companies, etc) and put it on webworm (password protected).

ANSS SM

CVTN optical isolation. Need to upgrade to year 3 hardware.

RDST needs upgrade to year 3 hardware with router.

NAIT: Couldn't get the NWS cable modem to cooperate at NAIT (firewall?). Dean was going to ask Comcast about it and maybe get a new modem. Steve will follow up.

Need to complete firmware upgrade on CVVA. The facility changed hands. They are willing to let us piggy back on their internet and John is working out the technical details with them (i.e. which side of the firewall we'll be on, hopefully the outside).

Now that CMG5 triggers actually contain data, we need to tune the trigger level. Mitch sent explanatory info and guidelines to Al and Steve to tune. Need to periodically test the data download function. Triggers need not be saved for any earthquakes less than magnitude 4 (we will already have data from nearly any event).

SLU had trouble getting triggers from the 10sps datastream. They are having success creating 200sps triggers off the 100sps channel without excessive triggers (though I wonder what their levels are). The two newer CMG5TD's we got from ASL also don't trigger well off the 10sps datastream. Jim will investigate the feasibility of switching to the 100sps tap as the trigger source for the 200sps upgrade. Then Al will retune as above.

Most of the strong motion stations can run at 100 sps without any modifications. We will install a router on the few that have problems (RDST will be the test case). These (eleven of them) will also need year 3 upgrades.

HDBR
Recorders 1, 6, and 7 are repaired and awaiting Steve to reconfigure. After that we can ask Bob for the next lane closure for reinstallation. Al reports that of the remaining four, one is not associating.

HDAR
The new battery controller is working. Need configure everything to the correct ip addresses, etc to get the data flowing to Boswell and on to CERI. Have a strong signal from HDBT. Need to determine what import machine and wsv all these data will be on. How do we trigger.

HDBT
Mud Island vertical array is operational. Not really, functional is a better word until we can get it online to qc/vet data (see HDAR).


ETSZ

Holly's ET list:
Down
TVNC (John has a replacement sensor)
SMNC (John has a replacement sensor)
CPRT (needs another solar panel)
GRBT- sporadic
CCRT- signal interference at certain times during the day

GMG.EHE - sensor needs replaced

Strong Motion Stations
C2SC
CVVA

STAR Node

RCGA charging circuitry at is old and primitive and in need of an upgrade.

GMG - Need battery box solar reg mounting plate and need to fuse the battery.

BRBC Node

GFM The NC state park service is purchasing GFM. We already have a couple permissions from them and John initiated the paperwork for the GFM permission.

TRYN: the forest service wants to remove the tower. There will be a new 180' tower and we have permission to be on it. We also have permission to be in the building (rack mount only). Can we convert this station to AC?

SWET:

CPCT: The aluminum enclosure with the GPS breakout board at CPCT is corroded, fragile, and needs replaced. Need to re-work lighting and pest chaser. John has a pest chaser and will use some landscape or pool hardware for the lighting. A new breakout and transient suppression enclosure is ready. Needs a locking tower box. This will also be upgraded to the standard bb.

FPAL The CMG3ESP is bad. John pulled it and is shipping it back to Memphis. There is still long period drift (period of about a day but not sinusoidal or monochromatic). The same behavior is present in other strongmotion stations (e.g. LNXT). Perhaps it is a thermal instability in the vault. This should be taken care of with an insulating sleeve on the broadband sensors.

GPS

James has a very nice SOH website at http://www.ceri.memphis.edu/people/jpdavis/gama/results.html.

Have permission to install freewaves at camp robinson so long as they don't interfere with anything currently operating (they shouldn't). Might be able to get on the base network. Need new permit.

Troy (HCES) took a severe lightning hit or massive power surge. James replaced a bunch of stuff and got it going again. Mattiolli sent an antenna and Bob loaned a reciever to get that colocated hardware going again.

Need to contact new owners (the water district) of the building at Miss. Co. Airport to refresh the permission. Chris contacted the county water dept will work with us any way they can. They also have internet on site there.

Need weed control and routine maintenance visit to all sites.

MAIR lost contact-access point probably needs rebooted.

Have a new ISP at Portageville. Now on DHCP but ISP is associating requests for old IP to correct mac address.

AG Network

All noise studies are complete. Horton is reviewing data to make siting choices. Gene Traverse from Nanometrics will be visiting Sepember 9. Mitch will invite Meremonte in for a visit. Mitch will check on Greg's status. Horton will persue permits.

Greg determined that current draw should be about 675mA at 12.0 VDC assuming a 70% transmit ratio for the freewave. Two 55-60 watt panels should suffice and to allow for temps down to 10 degrees F, we should use four 105ah batteries to get 14 days backup power.

FDL


MISCELLANEOUS

What to do with the Steiner brain dump DVD's?

Mitch made a seisnet group on umdrive. Go to http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password. There are several things on there including John's database. We can put contact info there too when its available. Only people in the seisnet group have access to this area. If you download something for update, don't forget to upload it again so that we all have the same version of the files.

The Oklahoma scream2scream connection doesn't seem to be as stable as we'd like so Steve is working with OGS to replace it with an ew export/import. The feed to NEIC is going but it would be good to verify with McManus at NEIC.

Rotational sensor is feeding an earthworm but still has IRIG crosstalk problem to fix. Its fine on an Oscope hooked to the output of the mux and looks fine in LabView. Its bad when viewed with wave_viewer. Need to see what the sac files look like to make sure it isn't just a wave_viewer feature.

Supply is checking the condition of Parker's old Suburban, Chris doubts it will run after a year sitting over there. He told her we would try to crank it up if needed. Looks like we could send old blue sometime soon. Maybe it would get on govdeals this year.

Pink Palace is working with Gary and Michelle on a new perc-like display area. We'll pull the receiver etc, when they start pulling out the old display (or sometime earlier if an opportunity arises).


CERI

Chris called P Plant electrical shop. They filled the generators soon after the outage. He said when there is a power failure as long as the generator is running that powers fuel tank pumps on campus they will bring fuel to keep us online. Chris will get all contact information posted near transfer switch and added to Jim's list.

3890 Power Chris is going to make a diagram for 3890. Good luck son.

3876 Power Chris is going to make a diagram for 3876. Good luck son.

Should we have a trouble ticket system?